A is three times as old as b. Four years ago A was four times as old as b was then. Find their present age.
Answer by macston(5194)

A=3B
A-4=4(B-4)
A-4=4B-16 Add 4 to each side.
A=4B-12 Substitute first equation for A
3B=4B-12 Subtract 3B from Each side
0=B-12 Add 12 to each side
12=B ANSWER 1: B is 12 years old
A=3B=3(12)=36 ANSWER 2: A is 36 years old
CHECK
A-4=4(B-4)
36-4=4(12-4)
32=4(8)
32=32